Posted by & filed under CSOs, EPA Updates, Frontpage, Remediation Work.

In-Canal Update

  • Supplemental dredging near the flushing tunnel to remove recently deposited debris is ongoing
    • Work began on May 9th
    • Work is anticipated to be completed by the end of July
  • Grouting of the armor mats in this area is being conducted concurrently as areas are dredged
  • Carroll Street bridge – no updates, 65% design for repair work is anticipated to be submitted in early summer
  • RTA 2 Construction
    • EPA reviewing access dredging work plan
    • 100% design for the dredging/capping/ISS work in RTA2 is expected to be submitted to EPA by mid-June

CSO Retention Tank Update

Red Hook Tank

  • Perimeter wall is approximately 80% complete.
    • 91 of 116 panels have been excavated and concreted as of May 24th (see attached figure)
    • Work is anticipated to be completed in July
  • Odor Complaints
    • In response to odor complaints from residents, primarily on Nevins and Sackett Streets, NYCDEP has taken the following steps over the past several months to minimize odors:
      • Continued operation of a perimeter misting system with odor masking scent, including a recent expansion of the system to the area south of the tank site (i.e., Sackett Street between Nevins Street and the canal)
      • Use of odor suppressing foam at the point of soil excavation, over on-site soil piles awaiting disposal, and, more recently, on soil as it is being loaded into trucks for off-site disposal.
      • Relocation of the soil loading area to a more centralized location on-site, away from the Nevins Street/Sackett Street intersection.
    • There is a Community Air Monitoring Plan in place which includes eight monitoring stations around the perimeter of the site.
      • Daily air monitoring data reports are compiled and posted on the NYCDEP website:
      • Exceedances of threshold levels in the CAMP have been infrequent, short duration, and often unrelated to excavation activities, such as fuel refilling operations or idling trucks near one of the monitoring stations.
    • NYCDEP and EPA oversight staff conduct multiple routine perimeter odor checks each day, as well as additional checks in response to any odor complaints.
    • NYCDEP has constructed a tent over certain equipment on-site which is believed to be a primary source of odors.
      • EPA has requested that the next phase of work, full-scale excavation of the RH-034 tank, be conducted completely within a tent structure.

Owl’s Head Tank

  • Demolition of buildings south of 5th Street began on May 21st
    • Expected to be completed by the end of June
  • Below ground tank construction work will begin with mobilization of equipment later this week.
    • Initial activities will include geotechnical and environmental borings beginning on June 10th.

Comments are closed.